You need to make an aqueous solution of 0.195 M barium bromide for an experiment in lab, using a 500 mL volumetric flask. How much solid barium bromide should you add?


1
Expert's answer
2022-02-28T11:01:15-0500

Number of moles= 0.195/2= 0.0975

Mass= rmm× moles

Mass= 297×0.0975= 29g